OpenStudy (anonymous):

sqrt(144) - sqrt(40) - sqrt(10) would equal Sqrt (94)? wouldn't it. or am i wrong?

OpenStudy (anonymous):

it would not

OpenStudy (anonymous):

Can you help me figure it out then please?

OpenStudy (anonymous):

\[\sqrt{a}+\sqrt{b}\neq \sqrt{a+b}\]

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

first off, what is the square root of 144 ?

OpenStudy (anonymous):

The equation thing isnt working it just looks like a bunch of random letters.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

refresh your browser and you should see it yes, the square root of 144 is 12, so we can start off with \[12-\sqrt{40}-\sqrt{10}\] can you see that? if now i will write it out without symbols

OpenStudy (anonymous):

crashed you still there?

OpenStudy (anonymous):

it works now.

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

haha.

OpenStudy (anonymous):

ok so now about \(\sqrt{40}\) we can write it as \[\sqrt{40}=\sqrt{4\times 10}=\sqrt{4}\sqrt{10}=2\sqrt{10}\]

OpenStudy (anonymous):

this gives you \[12-2\sqrt{10}-\sqrt{10}\] combine like terms, get \[12-3\sqrt{10}\]

OpenStudy (anonymous):

okay i got that now.

OpenStudy (anonymous):

let me know if you have any questions about any steps

OpenStudy (anonymous):

no I'm good thankyou.

OpenStudy (anonymous):

yw